Initially, there were equal amount of bananas and kiwis on the table. Sarah put 6 bananas and 8 kiwis in each basket. After the baskets were ready there were 16 bananas and 8 kiwis left on the table. How many baskets were made?
Let x be the number of baskets.
Equation :
Answer :
baskets were made.

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52786)

.

Equation:    6x + 16 = 8x + 8


Solution:    16 - 8  = 8x - 6x

             8       = 2x

             x        = 8/2 = 4.


ANSWER.  4 baskets.

Solved.


Answer by MathTherapy(10552)

Initially, there were equal amount of bananas and kiwis on the table. Sarah put 6 bananas and 8 kiwis in each basket. After the baskets were ready there were 16 bananas and 8 kiwis left on the table. How many baskets were made?
Let x be the number of baskets.
Equation :
Answer :
baskets were made.

There is a certain number of baskets
The number of bananas is the same as the number of kiwis
With number of baskets being x, we get: 6x + 16 = 8x + 8
6x - 8x = 8 - 16
- 2x = - 8
x, or